In aircraft flying at an altitude where the pressure was 35 kPa and

temperature – 38°C, stagnation pressure measured was 65.4 kPa. Calculate the speed of the

aircraft. Take molecular weight of air as 28


1
Expert's answer
2021-09-17T09:31:16-0400

The relation between a stagnation pressure and static pressure is the following:


"\frac{P_{t}}{P}=(1+\frac{\gamma-1}{2}M^2)^{\frac{\gamma}{\gamma-1}}";


Using the given data and the air "\gamma" coefficient we could rewrite the equation as:


"(1+0.2M^2)^{3.5}=1.869";


Now we could extract the Mach number M:


"M=\sqrt{5\cdot(1.869^{\frac{2}{7}}-1)}=0.989";


Finally, we could calculate the velocity of the aircraft:


"v=M\cdot c=0.989\cdot 343 m\cdot s^{-1}=339 m\cdot s^{-1}";


Thus, the aircraft is flying at 339 meters per second.